OpenStudy (anonymous):

The Product and Quotient Rule: Calculate dy dx You dont need to expand answer y = (5x^2+x)(x-x^2)

OpenStudy (anonymous):

how much can you do?

OpenStudy (anonymous):

not sure how to get the derivative -- im guessing the next step goes like this. 11(5x^2\[11(5x ^{2}+x)+(x-x ^{2})\]

OpenStudy (campbell_st):

the easiest way to do this is to expand the terms and get a polynomial y = -5x^4 + 4x^3 - x^2

OpenStudy (anonymous):

then what would be the next step to that?

OpenStudy (anonymous):

Product rule: \[\frac{\text{d}}{\text{d}x}f(x)g(x) = f'(x)g(x) + f(x)g'(x)\]So derivative first factor and left second one untouch, then derivative second one and left first one untouch.

OpenStudy (campbell_st):

now you can differentiate without the use of the product rule

OpenStudy (anonymous):

@campbell_st Would be better if s/he learn the product rule. Just saying.

OpenStudy (anonymous):

ok then how would i find the derivatives?

OpenStudy (anonymous):

what is f'(x), g'(x)?

OpenStudy (anonymous):

Derivative of the function. f(x) = (5x^2+x) and g(x) = (x-x^2) f'(x) = d/dx (5x^2+x) and g'(x) = d/dx (x-x^2) Does this make sense?

OpenStudy (campbell_st):

well the product rule can lead to mistakes... I think expanding makes it easier as most people are more confident with straight differentiation.

OpenStudy (anonymous):

so it would be (5x^2+x)(10)+(x-x^2)(2)?

OpenStudy (anonymous):

@campbell_st You have a point, but it actually depends on problem. If you face something like this: \[\Large \left(2x^3 + \frac{1}{\sqrt{x}}\right)^4 (3x^2 + x^{\frac{1}{4}})\] Do you think expanding this is easier than using product rule (and power rule)?

OpenStudy (anonymous):

No. First, derivative first factor (5x²+x), and left second one (x-x²) untouched. Do you know how to derivative (5x²+x)?

OpenStudy (campbell_st):

\[\frac{dy}{dx} = -20x^3 + 12x^2 - 2x\] is the derivative... so just answer the questions on there merits... for me... expand then differentiate will work. it can be written \[\frac{dy}{dx} = -2x(10x^2 - 6x + 1)\]

OpenStudy (anonymous):

@chrislb22 Do you want to use product rule or just do the campbell_st's way? This method does make it easier, but I think it'd be better for you to learn the product rule, just in case? It's up to you, though.

OpenStudy (anonymous):

product rule

OpenStudy (anonymous):

most likely i will be tested using the product rule, but in campbell's case it does make a bit more sense

OpenStudy (anonymous):

Well, just like I said, factor (5x²+x) and left (x-x²) untouched. Can you derivative (5x²+x)?

OpenStudy (anonymous):

derivative of (5x^2+x) is 11?

OpenStudy (anonymous):

10 from 5x^2 then 1 from x?

OpenStudy (anonymous):

Actually, while derivative the function, you basically subtract the power by one. So the derivative of 5x² is 10x. And, obviously, the derivative of x is one. So the derivative of 5x+x is 10x + 1.

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

oh i got you, i understand that now

OpenStudy (anonymous):

so the derivative for x-x^2 is 1-2x?

OpenStudy (anonymous):

Yeah, derivative first one and left second one untouched then "put in" plus sign then derivative second one and left first one untouched. Does this help?

OpenStudy (anonymous):

"so the derivative for x-x^2 is 1-2x?" Yes!

OpenStudy (anonymous):

Sorry, I accidentally deleted your message, what is your solution again?

OpenStudy (anonymous):

he Product and Quotient Rule: Calculate dy dx You dont need to expand answer y = (5x^2+x)(x-x^2)

OpenStudy (anonymous):

No, no, I mean your answer.